1. mass of aluminuum = 1.000g
  2. mass of beaker + dry alum (final product) = 39.00g
  3. mass of empty beaker = 25.000g
  4. mass of recovered dry alum (#2-#3) = 14.000g


Calculations:

  1. determining the limiting reagent: (listed reagents in the order used)

1.000 grams of Al = ________moles of Al

50 mL of 1.4 M KOH used = ________ moles of KOH

20 mL of 9.0 M H2SO4 = ___________moles of H2SO4


referring to the overall equation, the limiting reagent is? ____________


  1. theoretical yield calculation: _____ g ALUM
  2. Percent yield calculations: ________



  • list several common uses for alum?
  • explain how this process has a "green" application (hint: think recycling)
  • what are the hazards of this experiment? why? (list at least 3)
  • what errors would reduce the percent yield?



comments:

percent yield: 14.000/ theoretical yield x 100%


The theoretical yield is calculated using the overall equation:

2Al (s) + 2K+(aq) +2OH-(aq) + 8H+(aq) + 4SO42-(aq) + 22H2O(l) -> 2K[Al(SO4)2]. 12H2O + 3H2(g)

474.39 g/mol



1
Expert's answer
2021-03-03T01:33:14-0500

1. 1/27= 0,037037037 moles of Al.

50 x 1.4/1000 = 0,07 moles of KOH.

20 x 9/1000 = 0.18 moles of H2SO4.

Limiting reagent is Al.

0.07/2 =0.035.

0.035 x 27 = 0.945 g of Al.

Percent yield is 94.5%.

Uses of Alum


  • purification of drinking water as a chemical flocculant.
  • in styptic pencil to stop bleeding from minor cuts.
  • the adjuvant in vaccines ( a chemical that enhances the immune response)
  • deodorant "rock"
  • pickling agent to help keep pickles crisp.
  • flame retardant.
  • the acidic component of some types of baking powder.

Green chemistry is the design of chemical products and processes ... its design, manufacture, use, and ultimate disposal.

All hazards are assessed and categorized into three groups: biological, chemical and physical hazards. A general definition of a hazard as related to food safety is conditions or contaminants that can cause illness or injury.

Usually, percent yield is lower than 100% because the actual yield is often less than the theoretical value. Reasons for this can include incomplete or competing reactions and loss of sample during recovery.